• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fansler

    Log entry time 15:27:01 on March 16, 2012

    Entry number 366703

    keyword=LD1 Cryo/ Late Entry

    I went into the hall to check the LD1 He liquid level controls, and possibly swap a liquid level meter. The magnet He circuit has been experiencing problems since the 13th. I checked the probe length, and trip points, all were correct. I placed the meter in sample mode and it had no effect on the level reading. I checked the connectors on the back of the meter and they appeared to be intact. Ron Gilman texted me the Dipole dumped, so after lots of running around and head scratching I could find no reason for the fault except possibly a loose connector. Since the Dipole was ramped down, I swapped the meter, and inspected the connectors, one of the screws appeared to be cross threaded, problem solved.... but after several hours of monitoring, neither the connector, nor the new meter had any noticeable affect on the Dipole liquid level. The changes Matt made yesterday, seem to be correcting the problem, we will have to keep an eye on it over the weekend.

    A copy of this log entry has been emailed to: brads@jlab.org